OpenStudy (anonymous):

solving multi steps equatiotions r+1 ----= 8 3

OpenStudy (anonymous):

multiply both sides by 3 => (r+1)/3 X 3 = 8 X 3 => r+1 = 24 subtract 1 from both sides => r+1-1 = 24-1 => r = 23